# OpsEval: A Comprehensive IT Operations Benchmark Suite for Large Language Models

### v3

@@ 6.1. Implications of the evaluation results
  • Quantization with more than 3 bits can be an effective way to reduce computation and memory costs while preserving performance.
Overall, LLMs yield evaluation results on the OpsEval benchmark generally lower than those in general domains like MMLU (Hendrycks et al., 2021) and CEval (Huang et al., 2023). This comparison highlights the necessity of explicitly fine-tuning OpsLLM for the Ops field.
Our tests on the model with various prompt engineering techniques indicate that prompt engineering significantly impacts the effectiveness of eliciting operations knowledge from the model. This suggests that further research into prompt engineering is needed to enhance the performance of large models in this specific vertical domain.
6.2. Automated QA generation for question-answering questions
During the data collection process, we have considered automating question-answer generation. First, we sampled the question-answer pairs and manually assessed their accuracy and domain relevance. Later, we used typical manual evaluation examples for few-shot learning, enabling GPT to evaluate question-answer pairs based on our evaluation criteria automatically.
Directly generated question-answers tend to be simple judgment or concept questions rather than reasoning questions that better demonstrate the model’s capabilities and knowledge density. Our goal is to ensure that while the topics of the questions remain relevant to the seed questions, their specific content is distinct from the original questions. By maintaining the overarching framework in the Ops domain, we can expand the number and types of questions, enabling a more comprehensive evaluation of model capabilities. Additionally, we can incorporate external knowledge during the data generation, continually enhancing our ability to evaluate new content.
6.3. Examination Methodology of Hallucination Issues in LLM within OpsEval
Recognizing the necessity for LLMs to alleviate the hallucination problem, we explore the corresponding evaluation methodology.
Our approach involves both multi-choice and question-answering evaluations. Multi-choice questions assess the LLMs based on their knowledge and diverse Ops capabilities. However, to effectively uncover hallucinations, we rely on question-answering evaluation. This involves expert review, where outputs are scrutinized against three main criteria: Fluency, Accuracy, and Evidence. The Evidence criterion, in particular, is pivotal in highlighting hallucination issues. It underscores the significance of LLMs providing evidence or a reasoning process to support their outputs.
The intersection of hallucination problems and evidence-based evaluation is a focal point in our research. By emphasizing evidence, we aim to mitigate the issue of hallucinations, ensuring that LLM outputs are not only plausible but also verifiable. We plan to refine our multi-choice question design. This will enable us to automate the evaluation of model robustness against hallucinations more effectively. This adjustment is anticipated to enhance further our understanding of LLMs’ capabilities and limitations in dealing with hallucinations, thereby contributing to developing more reliable and trustworthy LLMs in the Ops field.
7. Conclusion
In this paper, we introduce OpsEval, a comprehensive Ops benchmark suite designed for LLMs. Unprecedented in its approach, OpsEval evaluates LLMs’ proficiency across various Ops scenarios while considering varying ability levels encompassing knowledge recall, analytical thinking, and practical application. This comprehensive benchmark comprises 7,200 questions in both multi-choice and question-answering formats, both English and Chinese. After careful review by experts from the industry, we have made the datasets in OpsEval available to the public on Github, and publish the evaluation leaderboard on the official OpsEval website.
Supported by quantitative and qualitative results, we elucidate the nuanced impact of various LLM techniques, such as chain-of-thought and few-shot in-context learning, on Ops performance. Notably, the GPT4-score emerges as a more reliable metric when compared to widely used BLEU and ROUGE, suggesting its potential as a replacement for manual labeling in large-scale quantitative evaluations.
@@ 7. Conclusion
The identified flexibility within the OpsEval framework presents opportunities for future exploration. The adaptability of this benchmark facilitates the seamless integration of additional fine-grained tasks, providing a foundation for continued research and optimization of LLMs tailored for Ops.
@@ Appendix A Details of OpsEval Benchmark
A.1. Information on the companies and experts participating in OpsEval
Table 10. Information of companies collaborating in OpsEval
@@ A.1. Information on the companies and experts participating in OpsEval
A.2. Task Types of Questions
@@
A.3. Ability Levels of Questions
@@
Figure 7. Three examples of the processed questions
@@
A.4. An Example of Subjective Questions
Figure 8. An example of the saved subjective questions
@@ A.4. An Example of Subjective Questions
Figure 9. An example of building the prompt of subjective questions.
@@
Table 11. LLMs’ overall performance on 5G communication operations test set
@@
Note: The best accuracy of each language for each LLM is in bold font. The best accuracy of all LLMs for each setting is underlined.
Table 12. LLMs’ overall performance on database operations test set
@@
Qwen-14B-Chat 24.95 28.37 33.00 36.62 27.97 28.37 27.97 24.14 27.57 27.57 32.39 36.02 40.04 35.41 30.38 33.4
DevOps-Model-14B-Chat 25.15 26.96 35.41 38.83 33.2 34.81 27.36 27.36 24.75 22.74 28.37 27.77 36.62 37.02 27.57 26.36
LLaMA-2-70B 19.72 19.72 27.97 27.97 26.56 26.56 32.6 32.6 15.29 15.29 34.81 34.81 26.76 26.76 33.8 33.8
@@
Note: The best accuracy of each language for each LLM is in bold font. The best accuracy of all LLMs for each setting is underlined.
A.5. Metrics used in Question-Answering Evaluation
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) is a set of metrics for evaluating machine translation and summarization. ROUGE-N is the overlap of n-grams between the prediction and reference. ROUGE-L assesses sentence structure similarity and identifies the longest sequential n-grams. ROUGE can be understood as the recall of the ground-truth answer. The score of ROUGE is normalized from 0 to 100. The higher the score is, the better it is.
BLEU (Bilingual Evaluation Understudy) can be understood as the precision of the generated answer. We utilize the scarebleu python package to calculate BLEU in OpsEval. The score of BLEU is normalized from 0 to 100. The higher the score is, the better it is.
GPT4-Score is a score generated by GPT4 with a deliberately crafted prompt. Scoring by LLMs is used increasingly (Bai et al., 2023) (Chang et al., 2023), especially after the parameters of LLMs get larger. We compose the scoring prompt of the question, the ground-truth keypoint, the ground-truth detailed answer, and the answer of LLM to be scored. The score is between 1 and 10, and the higher is better. The prompt for GPT-4 Scoring is shown in Figure 2(b).
Appendix B Additional details of experiments
B.1. Detailed Information of LLMs Evaluated
GPT-4 (OpenAI, 2023a) is a large multimodal model (accepting image and text inputs, emitting text outputs) that, while less capable than humans in many real-world scenarios, exhibits human-level performance on various professional and academic benchmarks. It is recognized as the strongest lanuage model currently. ChatGPT (OpenAI, 2022) is an earlier AI-powered language model developed by OpenAI which is built upon GPT-3.5. We use the GPT-3.5-turbo version in our experiments. LLama 2 (et.al., 2023) is a second-generation open-source LLM from Meta which is very popular due to its open-source feature. It has the ability to process multiple languages including Chinese. We evaluate three weights (70B, 13B and 7B as shown in 3) of LLama 2.
Although LLama 2 is able to process Chinese input, it has a small Chinese vocabulary so that its abitilty of understanding and generating Chinese text is limited. As a result, we evaluate some Chinese-oriented LLMs which are published by institutions in China. ERNIE-Bot 4.0 (202, 2024a) is the latest self-developed language model released by Baidu. As claimed by Baidu, ERNIE-Bot 4.0 rivals OpenAI’s GPT-4. Qwen (202, 2023) (abbr. Tongyi Qianwen) is a series of LLMs developed by Alibaba Cloud. And Qwen-Chat is a series of large-model-based AI assistant trained with alignment techniques based on the pretrained Qwen. We evaluate three weights (72B, 14B and 7B as shown in 3) of Qwen-Chat. Baichuan2-13B-Chat (Baichuan, 2023) is aligned chat model based on Baichuan2-13B-Base (Baichuan, 2023) which is an open-source LLM published by Baichuan Intelligence. GLM (Du et al., 2022), developed by Tsinghua Knowledge Engineering Group, is a General Language Model pretrained with an autoregressive blank-filling objective and can be finetuned on various natural language understanding and generation tasks. Based on GLM, Zhipu AI released GLM4 (the newest version of GLM model) (202, 2024b) and GLM3 (the third version of GLM model). For GLM3, we use GLM3-turbo (202, 2024b) version and ChatGLM3-6B (202, 2024b) in our experiments. InternLM2-Chat-20B and InternLM2-Chat-7B (Team, 2023), recently developed by Shanghai AI Laboratory, are multi-lingual models based on billions of parameters through multi-stage progressive training on over trillions of tokens. Furthermore, we evaluate DevOps-Model-14B-Chat (devopspal2023), an open source Chinese DevOps oriented models, mainly dedicated to exerting practical value in the field of DevOps.
B.2. An Example of CoT Prompt
Figure 10. An example of zero-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
Figure 11. An example of few-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
B.3. Overview Performance on Different Test Sets
@@
In Table 11 and Table 12, we present overview performance of different LLMs on the 2 test sets in OpsEval, including 5G Communication Technology Operations and Database Operations.
B.4. Performance on Different Quantization Models
Figure 12 shows the accuracy of LLaMA-2-70B of different quantization parameters on Chinsese objective questions. We do both few-shot and zero-shot evaluation with the naive setting.

### v4

@@ 4.4 Performance on Different Quantization parameters
Practical Lesson: Quantization with more than 3 bits can effectively reduce computation and memory costs while preserving performance.
5 Validation
5.1 Benchmark Leakage Test
Table 6: Measurement of potential test data leakage during the training of LLM. This demonstrates the unbiased nature and non-leakage of the OpsEval test set.
Dataset $L_{test}$ $L_{ref}$ $\Delta L$
Alpaca 1.994033 2.354260 -0.360228
Alpaca-GPT4 1.498862 1.763663 -0.391062
CEval 2.570809 2.309943 0.260866
MMLU 2.547598 2.189870 0.357728
OpsEval 1.885437 1.728079 0.105095
For the fairness of a benchmark suited for LLM, avoiding potential bias emerging from test set leakage is necessary. We adapted the methodology from Wei, et.al. (2023) to perform a leakage test on OpsEval’s dataset. We evaluate the LLM loss on samples from different datasets for several LLMs and calculate the average loss. For each dataset, we compare LLM loss on the test split ($L_{test}$) and a specially curated reference set ($L_{ref}$ ) generated by GPT-4, designed to mimic the testing dataset. While  Wei, et.al. (2023) only asked GPT-4 to generate similar questions to the GSM8K (Cobbe et al., 2021) dataset, we require GPT-4 to rewrite the question while preserving its original meaning and accuracy. We define a key metric: $\Delta L=L_{test}-L_{ref}$, serving as an indicator of potential test data leakage. A lower $\Delta L$ suggests that the LLM’s lower $L_{test}$ comes from overfitting the test set rather than understanding the questions, indicating potential leakage. Table 6 shows the results of leakage measurement. In addition to the two standard evaluation benchmarks (CEval (Huang et al., 2023) and MMLU (Hendrycks et al., 2021)), we conducted the same experiments on the alpaca dataset (Taori et al., 2023) and the Alpaca-GPT4 dataset (Peng et al., 2023), which is likely used in the pre-training of large models, using its $\Delta L$ as reference. The corpora likely involved in training show a significantly smaller $\Delta L$, whereas the loss for the OpsEval dataset remains at a relatively small positive value. This demonstrates the unbiased nature and non-leakage of the OpsEval test set. The models we used in the leakage test are listed in Appendix B.1.
5.2 Expert alignment of FAE-Score
Table 7: Pearson correlation coefficients between Expert-Evaluation metrics and Automated metrics. Total is the sum of Fluency, Accuracy, and Evidence.
Metric Total Fluency Accuracy Evidence
FAE-Score 0.9175 0.7200 0.9799 0.7962
BLEU-Score 0.6705 0.8253 0.6004 0.4281
ROUGELsum -0.3957 -0.2893 -0.0814 -0.6660
Table 7 shows the correlation coefficients between various automated scoring metrics (ROUGE, BLEU, and FAE-Score) and Expert-Evaluation sub-metrics. The results indicate that ROUGE and BLEU scores often misalign with Expert-Evaluation. This misalignment occurs because LLMs with poor performance may generate keywords that boost ROUGE and BLEU scores, while stronger LLMs might receive lower scores due to different wording from standard answers. In contrast, FAE-Score rankings closely match Expert-Evaluation, particularly with the Accuracy metric. This suggests that FAE-Score is more reliable in assessing the factual accuracy of LLMs’ outputs. Notably, GPT-4’s performance in factual accuracy is reflected in its strong alignment with the Accuracy metric.
6 Limitation and Future Work
Despite the positive contributions of this study, we recognize the following limitations: 1) Topic Distribution Imbalance: There may be an uneven distribution of topic classifications. This issue can be addressed by consciously supplementing with community contributions in future iterations. 2) Agent and RAG Introduction: The inclusion of agents and Retrieval-Augmented Generation (RAG) techniques is constrained by the current large models’ lack of foundational knowledge in operations. Our leaderboard will incorporate more complex tasks once open-source models possess sufficient operational capabilities. 3) Reproducibility of FAE-Score: FAE-Score requires a certain level of instruction-following ability from the judge model, which is why we chose GPT-4 for this work. However, it is best to use open-source models as judge model to ensure reproducibility. In our subsequent work, we will also use open-source, smaller models like Qwen2-72B for evaluation. We believe that any model that ranks higher than Qwen2-72B on general benchmarks, such as SuperGLUE[1], can be considered a suitable judge model. 4) Potential Negative Societal Impact: The use of private domain data from companies necessitates strict adherence to data usage permissions to avoid potential privacy breaches.
7 Conclusion
In this paper, we introduced OpsEval, the first comprehensive Ops benchmark suite designed for evaluating the performance of large language models (LLMs) in IT operations. We established a robust evaluation framework encompassing a wide range of sub-domains and tasks within Ops through rigorous data collection from multiple sources and meticulous preprocessing steps. Our benchmark includes a carefully selected set of 8920 questions, which we have partially released to aid initial evaluations while protecting the integrity of the remaining dataset. It has undergone experiments in data leakage detection, ensuring its reliability. Our observations, supported by quantitative and qualitative results, highlight the need for a balanced approach to selecting fundamental models, considering both performance and robustness. During the QA evaluation, the FAE-Score emerges as a more reliable metric than traditional metrics, suggesting its potential as a replacement for manual labeling in large-scale quantitative evaluations. Our failure rate analysis across 8 tasks and 3 abilities provides researchers with crucial insights and prospects for future breakthroughs.
@@ 7 Conclusion
The identified flexibility within the OpsEval framework presents opportunities for future exploration. This benchmark’s adaptability facilitates the seamless integration of additional fine-grained tasks, providing a foundation for continued research and optimization of LLMs tailored for Ops.
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China under Grant 62072264, and the Beijing National Research Center for Information Science and Technology (BNRist) key projects.
@@ Appendix A Details of OpsEval Benchmark
A.1 Information on the companies and experts participating in OpsEval
Table 8: Information of companies collaborating in OpsEval
@@ A.1 Information on the companies and experts participating in OpsEval
A.2 Dependance Filtering Keyword List
question_keywords = [‘the figure’, ‘the scenario’, ‘the previous question’] fail_pred_keywords = [‘unclear’, ‘scenario is not provided’, ‘cannot be determined’, ‘none of the options’, ‘none of the given options’]
A.3 Task Types of Questions
@@
A.4 Ability Levels of Questions
@@
Figure 6: Three examples of the processed questions
@@
A.5 Prompt and Formatting of Questions
Figure 7: Three examples of the processed questions
A.6 An Example of Subjective Questions
Figure 8: An example of the saved subjective questions
@@ A.6 An Example of Subjective Questions
Figure 9: An example of building the prompt of subjective questions.
@@
A.7 Automated QA generation
During the data collection process, we have experimented automating question-answer generation. We first sampled the QA pairs and manually assessed their accuracy and domain relevance. Later, we used typical manual evaluation examples for few-shot learning, enabling GPT to evaluate QA pairs based on our evaluation criteria automatically. Directly generated question-answers tend to be simple judgment or concept questions rather than reasoning questions that better demonstrate the model’s capabilities and knowledge density. Our goal is to ensure that while the topics of the questions remain relevant to the seed questions, their specific content is distinct from the original questions. By maintaining the overarching framework in the Ops domain, we can expand the number and types of questions, enabling a more comprehensive evaluation of model capabilities. Additionally, we can incorporate external knowledge during the data generation, continually enhancing our ability to evaluate new content.
Appendix B Additional details of experiments
B.1 Detailed Information of LLMs Evaluated
Table 9: Models evaluated in this paper. The “access” column in the table shows whether we have full access to the model weights or can only access them through API.
Model Creator #Parameters Access License
GPT-4/GPT-3.5-turbo OpenAI undisclosed API Proprietary
ERNIE-Bot-4.0 Baidu undisclosed API Proprietary
GLM4/GLM3-turbo Tsinghua Zhipu undisclosed API Proprietary
Meta-LLaMA-3 Meta 8B Weights Llama 3 Community
LLaMA-2 Meta 7/13/70B Weights Llama 2 Community
Qwen-Chat Alibaba Cloud 7/14/72B Weights Qianwen LICENSE
Qwen1.5-Chat Alibaba Cloud 14B Weights Qianwen LICENSE
InternLM2-Chat Shanghai AI Laboratory 7/20B Weights Apache-2.0
DevOps-Model-Chat CodeFuse 14B Weights Apache-2.0
Baichuan2-Chat Baichuan Intelligence 13B Weights Apache-2.0
ChatGLM3 Tsinghua Zhipu 6B Weights Apache-2.0
Mistral Mistral 7B Weights Apache-2.0
Gemma Google 2/7B Weights Gemma license
GPT-4 (OpenAI, 2023a) is a large multimodal model (accepting image and text inputs, emitting text outputs) that, while less capable than humans in many real-world scenarios, exhibits human-level performance on various professional and academic benchmarks. It is recognized as the strongest lanuage model currently. ChatGPT (OpenAI, 2022) is an earlier AI-powered language model developed by OpenAI which is built upon GPT-3.5. We use the GPT-3.5-turbo version in our experiments. LLaMA 2 (Touvron, et.al., 2023) is a second-generation open-source LLM from Meta which is very popular due to its open-source feature. It has the ability to process multiple languages including Chinese. We evaluate three weights (70B, 13B and 7B as shown in 9) of LLaMA 2.
Although LLaMA 2 is able to process Chinese input, it has a small Chinese vocabulary so that its abitilty of understanding and generating Chinese text is limited. As a result, we evaluate some Chinese-oriented LLMs which are published by institutions in China. ERNIE-Bot 4.0 (202, 2024) is the latest self-developed language model released by Baidu. As claimed by Baidu, ERNIE-Bot 4.0 rivals OpenAI’s GPT-4. Qwen (202, 2023) (abbr. Tongyi Qianwen) is a series of LLMs developed by Alibaba Cloud. And Qwen-Chat is a series of large-model-based AI assistant trained with alignment techniques based on the pretrained Qwen. We evaluate three weights (72B, 14B and 7B as shown in 9) of Qwen-Chat. Baichuan2-13B-Chat (Baichuan, 2023) is aligned chat model based on Baichuan2-13B-Base (Baichuan, 2023) which is an open-source LLM published by Baichuan Intelligence. GLM (Du et al., 2022), developed by Tsinghua Knowledge Engineering Group, is a General Language Model pretrained with an autoregressive blank-filling objective and can be finetuned on various natural language understanding and generation tasks. Based on GLM, Zhipu AI released GLM4 (the newest version of GLM model) (Zeng et al., 2022) and GLM3 (the third version of GLM model). For GLM3, we use GLM3-turbo (Zeng et al., 2022) version and ChatGLM3-6B (Zeng et al., 2022) in our experiments. InternLM2-Chat-20B and InternLM2-Chat-7B (InternLM_Team, 2023), recently developed by Shanghai AI Laboratory, are multi-lingual models based on billions of parameters through multi-stage progressive training on over trillions of tokens. Furthermore, we evaluate DevOps-Model-14B-Chat (AI, 2024a), an open source Chinese DevOps oriented models, mainly dedicated to exerting practical value in the field of DevOps.Gemma  (Gemma_Team et al., 2024) is a family of lightweight, state-of-the-art open models based on Gemini technology from Google DeepMind. Trained on up to 6T tokens, Gemma achieves excellent language understanding and reasoning capabilities. We conducted an evaluation of Gemma-2b and Gemma-7b to investigate the effectiveness of Gemma with different weights.
Table 10: GPTQ models for LLaMA-2-70B
Model Size #GPTQ Dataset Disc
LLaMA-2-70B 140GB / Raw LLaMA-2-70B model.
LLaMA-2-70B-Int4 35.33GB wikitext 4-bit quantization model.
LLaMA-2-70B-Int3 26.78GB wikitext 3-bit quantization model.
B.2 Prompts
Figure 10: An example of zero-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
Figure 11: An example of few-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
B.3 Compute and Resources Used for Experiments
During our OpEval experiments evaluating different LLMs, we utilize an 8 Nvidia A800-80GB GPU cluster to run inference on models with available weights. For models with API access, we perform inference using CPUs.
B.4 Overview Performance on Different Test Sets
@@
Table 11: LLMs’ overall performance on wired network operations test set
@@
GPT-4 / / / / / / 88.70 / / / / / / / 86.00 /
Qwen-72B-Chat 70.41 70.50 72.38 72.56 70.32 70.32 70.13 70.22 65.77 65.86 68.13 68.30 69.40 69.40 69.99 70.08
GPT-3.5-turbo 66.60 66.80 69.60 72.00 68.30 68.30 70.90 72.50 58.40 58.60 64.80 67.60 59.20 59.70 65.20 67.40
ERNIE-Bot-4.0 61.15 61.15 70.00 70.00 60.00 60.00 70.00 70.00 67.54 67.54 71.96 71.96 72.00 72.00 78.00 78.00
Qwen1.5-14B-Chat 54.90 34.88 64.09 60.82 52.23 65.55 59.54 47.08 54.04 45.18 62.56 59.12 58.78 61.10 63.43 52.5
Devops-Model-14B-Chat 30.69 30.59 55.77 63.63 63.85 61.96 41.15 44.01 47.59 46.57 52.52 56.01 62.07 60.08 50.59 55.79
Qwen-14B-Chat 43.78 47.81 56.58 59.40 62.09 59.70 49.06 55.88 48.35 48.81 55.35 57.40 58.53 56.12 52.12 54.99
LLaMA-2-13B 41.80 46.50 53.10 58.70 53.30 53.00 56.80 61.00 29.70 31.60 51.60 57.00 39.60 38.90 48.00 50.60
Gemma-7B 25.09 25.09 50.86 50.86 59.12 59.12 50.77 50.77 31.58 31.58 47.59 47.59 34.68 34.68 48.88 48.88
LLaMA-2-70B-Chat 25.29 25.29 57.97 58.06 52.97 52.97 58.55 58.55 38.55 38.55 57.49 57.49 49.09 49.09 48.57 48.57
Internlm2-Chat-20B 56.36 56.36 26.18 26.18 60.48 60.48 45.10 45.10 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
Internlm2-Chat-7B 49.74 49.74 56.19 56.19 48.20 48.20 49.74 49.74 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
LLaMA-2-7B 39.50 40.00 45.40 49.50 48.20 46.80 52.00 55.20 29.80 30.20 50.10 55.60 38.60 40.80 45.60 50.40
Qwen-7B-Chat 45.90 46.00 47.30 50.10 52.10 51.00 48.30 49.80 29.60 29.90 50.60 53.50 50.40 46.90 46.90 47.70
Baichuan2-13B-Chat 37.90 38.30 42.70 46.60 51.90 51.60 44.50 47.45 44.60 45.40 41.60 44.30 45.60 45.70 43.90 46.70
Note: The best accuracy of each language for each LLM is in bold font.
Table 12: LLMs’ overall performance on 5G communication operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
GPT-4 / / 56.30 65.49 / / 59.62 63.54 / / 57.19 62.11 / / 61.55 65.68
@@
Gemma-2B 20.10 20.10 24.20 24.20 31.20 31.20 35.50 35.50 25.60 25.60 28.30 28.30 19.10 19.10 35.50 35.50
Note: The best accuracy of each language for each LLM is in bold font.
Table 13: LLMs’ overall performance on database operations test set
@@
Qwen-14B-Chat 24.95 28.37 33.00 36.62 27.97 28.37 27.97 24.14 27.57 27.57 32.39 36.02 40.04 35.41 30.38 33.40
DevOps-Model-14B-Chat 25.15 26.96 35.41 38.83 33.20 34.81 27.36 27.36 24.75 22.74 28.37 27.77 36.62 37.02 27.57 26.36
LLaMA-2-70B 19.72 19.72 27.97 27.97 26.56 26.56 32.6 32.6 15.29 15.29 34.81 34.81 26.76 26.76 33.80 33.80
@@
Gemma-2B 16.90 16.90 19.52 19.52 16.10 16.10 24.75 24.75 18.51 18.51 24.95 24.95 21.53 21.53 27.77 27.77
Note: The best accuracy of each language for each LLM is in bold font.
@@
In Table 11, Table 12 and Table 13, we present overview performance of different LLMs on the 3 test sets in OpsEval, including Wired Network Operations, 5G Communication Technology Operations and Database Operations.
B.5 Performance on Different Quantization Models